OpenStudy (anonymous):

how do you find the supplementary and complementary angle of 17˚

myininaya (myininaya):

sup=x 180=x+17 com=y 90=y+17

OpenStudy (anonymous):

is it for a triangle or circle supplement would be 180 - 17 for triangle or 360-17 for circle

OpenStudy (anonymous):

sup is 180 comp is 90 subtract 17 from both for your answers

OpenStudy (anonymous):

supplementary angle of 17* is 180*-17*=163* and complementary angle of 17* is 90*-17*=73*

OpenStudy (anonymous):

Tank you all!!(: